### v1

# Manolis Mylonas CERTH-ITI Thessaloniki, Greece

Abstract
In this work, we introduce the task of script-driven video summarization, which aims to produce a summary of the full-length video by selecting the parts that are most relevant to a user-provided script outlining the visual content of the desired summary. Following, we extend a recently-introduced large-scale dataset for generic video summarization (VideoXum) by producing natural language descriptions of the different human-annotated summaries that are available per video. In this way we make it compatible with the introduced task, since the available triplets of “video, summary and summary description” can be used for training a method that is able to produce different summaries for a given video, driven by the provided script about the content of each summary. Finally, we develop a new network architecture for script-driven video summarization (SD-VSum), that relies on the use of a cross-modal attention mechanism for aligning and fusing information from the visual and text modalities. Our experimental evaluations demonstrate the advanced performance of SD-VSum against state-of-the-art approaches for query-driven and generic (unimodal and multimodal) summarization from the literature, and document its capacity to produce video summaries that are adapted to each user’s needs about their content.
Keywords Script-driven video summarization  $\cdot$ Dataset  $\cdot$ Natural language descriptions  $\cdot$ Cross-modal attention  $\cdot$ Large Multimodal Models
1 Introduction
The established approach in the Media industry for producing summarized versions of a full-length video for distribution via different communication channels (e.g., social networks and video sharing platforms), requires the visual inspection of the entire content by a professional editor who then decides about the parts that should be included in the summary. This time-demanding and laborious task (especially in the case of long videos with diverse visual content) can be facilitated by modern AI-based video summarization technologies [1]. Nevertheless, these technologies can assist the work of editors only to a certain extent, since they typically learn how to create video summaries based on general-purpose summarization criteria, such as the representativeness and diversity of the summary (see the top part of Fig. 1), and thus, are not capable of generating video summaries according to specific needs of the editor about their content.
Figure 1: A video blog (vlog) with the daily activities of the vlogger, and the produced summaries by: generic video summarization (top row), query-driven video summarization (middle row), and script-based video summarization (bottom row). The generic video summary aims to provide a complete overview of the video. The query-driven video summary concentrates solely on parts of the video that match with a short user query. The script-driven video summary offers a more diverse and complete synopsis of the video, containing parts that relate to the different activities outlined in the user script, and avoiding parts the are less interesting for the user, and thus redundant.
Being more tailored to the aforementioned need, a number of text/query-driven video summarization methods have been described in the literature [2, 3, 4, 5, 6, 7, 8]. These methods allow to generate video summaries that are more customized to the users’ preferences. However, these preferences are typically expressed using one or more keywords [2, 3, 6] or a short sentence [4, 5]. Hence, the existing methods for text/query-driven video summarization are not compatible with more extensive descriptions about the content of the summary. As a consequence, the generated summaries by these methods exhibit limited visual and semantic diversity, as they mainly contain parts of the video that match a short user query (see the middle part of Fig. 1).
To overcome this limitation, in this paper we introduce the task of script-driven video summarization, whose goal is to generate a summary of the full-length video by choosing the parts which are most relevant to a longer user-provided query (“script”) that outlines the visual content of the desired summary. To support research on this task, we extend a recently-introduced large-scale dataset for video summarization (VideoXum [9]), by producing natural language descriptions of the visual content of the different human-annotated summaries that are available per video, using the state-of-the-art Large Multimodal Model LLaVA-NeXT-7B [10]. The multiple triplets of “video, summary and summary description” that are available for each video in the extended S-VideoXum dataset can be used to train a method to produce different summaries for a given video, based on the provided description (the so-called script) about the content of each summary. Finally, we build a new method for script-driven video summarization (called SD-VSum), that employs a cross-modal attention mechanism for aligning and fusing information from the visual and text modalities, and computes the frames’ importance using a trainable Transformer-based scoring mechanism. SD-VSum is designed to generate video summaries that are visually and semantically more diverse (see the bottom part of Fig. 1) compared to summaries created by text/query-driven summarization methods, and less redundant compared to summaries produced by generic summarization approaches. We train and evaluate SD-VSum using S-VideoXum and compare its performance against state-of-the-art approaches for query-driven and generic (unimodal and multimodal) video summarization from the literature. Our quantitative and qualitative evaluations highlight the advanced performance of SD-VSum and demonstrate its capacity to generate video summaries that meet the users’ needs. Our contributions are as follows:
  • We introduce the task of script-driven video summarization, which aims to produce a summary of the full-length video by selecting the parts that are most relevant to a user-provided script about the visual content of the summary.
  • We extend the recently-introduced VideoXum large-scale dataset for video summarization to S-VideoXum, by augmenting the former with natural language descriptions of the different human-annotated summaries that are available per video.
  • We propose a new method for script-driven video summarization (SD-VSum) that employs a cross-modal attention mechanism; we use the extended S-VideoXum dataset to train and evaluate this method, and document its competitiveness against other state-of-the-art for query-driven and generic (unimodal and multimodal) video summarization.
2 Related Work
2.1 Text/Query-driven video summarization
In one of the first attempts towards text-driven video summarization, Sharghi et al. (2016) [2] proposed the use of a probabilistic model for query-focused video summarization, which creates a summary by selecting the video shots that jointly have the highest relevance to the user query and the highest importance in the context of the video. To avoid the use of costly supervision data, in their subsequent work, Sharghi et al. (2017) [3] described a memory network parameterized probabilistic model that implicitly attends the user query about the video onto different video frames and shots. Vasudevan et al. (2017) [11] utilized the framework of submodular optimization to form a method for query-adaptive video summarization that selects a subset of frames for inclusion in the video summary by quantifying their relevance with the text query based on the distance of deep visual and text representations in a common embedding space, and taking into account the frames’ diversity, representativeness and aesthetic quality. Wei et al. (2018) [4] presented a semantic attended video summarization network that contains a frame selector and a video descriptor and learns how to choose the most semantically representative video parts by minimizing the distance between the generated description for the video summary, and human-provided descriptions of visually-coherent segments of the video. Zhang et al. (2018) [12] suggested the use of a query-conditioned generative adversarial network, where the generator learns the joint representation of the user query and the video content, the discriminator tries to discriminate the real summary from a generated and a random one, and a three-player loss is used to force the generator to learn the task and avoid random trivial solutions. Jiang et al. (2019) [13] designed a hierarchical network architecture to model query-related long-range temporal dependency and used diverse attention mechanisms to encode query-related and context-important information in combination with a multi-level self-attention and a variational autoencoder to incorporate user-oriented diversity and stochastic factors. Huang et al. (2020) [5] described a method for query-controllable video summarization, that consists of: a summary controller that uses text (e.g., words, phrases, sentences) to describe the desired video summary, a summary generator that scores the video frames based on their implicit relationship with the text-based description, and a summary output module that creates a video summary by selecting the top-k scoring frames. Xiao et al. (2020) [14] proposed the use of a fully convolutional network architecture which encodes the high-level semantics of each video shot using local self-attention mechanisms and models the semantic relationship between all shots and a given query using a query-aware global-attention, in combination with a query-relevance ranking module which computes a similarity score for each shot for a given query and forms the summary. Aiming to eliminate the bias when the model is trained on a small dataset, in their next work, Xiao et al. (2020) [15] trained a hierarchical self-attentive network for estimating the importance of video frames and shots using the ActivityNet Captions dataset [16], fine-tuned it using a reinforced caption generator that aims to construct the video description using the selected frames, and employed a query-aware scoring module that computes shot-level scores for a given query and produces the query-related video summary. Cizmeciler et al. (2021) [17] presented a weakly-supervised query-focused video summarization method that makes use of semantic attributes as an indicator of query relevance and semantic saliency maps obtained from predictions of pretrained attribute/action classifiers to locate relevant regions in the video frames, and employs both within a submodular maximization framework. Narasimhan et al. (2021) [6] introduced the CLIP-It method for query-focused video summarization, which relies on a multi-head language-guided attention mechanism that learns how to estimate the frames importance based on their visual relevance and their correlation with a user-defined query. Formulating the task of query-attentive video summarization as a multi-label classification problem, Hu et al. (2023) [18] predicted the correlation between visual content and multi-concept labels by feeding convolutional features into a multi-layer perceptron, used the labels’ cross-correlation to weight the predicted probability, and selected the video part with the highest relevance to the user’s query sentence to form the summary. Finally, to deal with the lack of large labeled datasets for query-based video summarization, Huang et al. (2023) [19] resorted to the use of self-supervision and the generation of segment-level pseudo labels from input videos to model the relationship between a pretext and a target task and between the pseudo label and the human-defined label, and employed a semantics booster to generate context-aware query representations and mutual attention to capture the interaction between visual and textual modalities. For further details, we refer the reader to a recently published survey on query-attentive video summarization [20].
The methods described above create a video summary taking into account the user’s preferences about its content, expressed using a few keywords [2, 3, 11, 12, 13, 5, 14, 17, 15, 6, 18, 19] or a short sentence [4, 6]. Contrary to these methods, SD-VSum guides the summarization process using a free-form language description of the content of the desired video summary, allowing the user to create summaries that are visually and semantically more diverse compared to the summaries generated by the existing text/query-driven summarization methods.
2.2 Multimodal video summarization
Closely related to the text/query-driven videos summarization methods, in terms of the used data modalities, are multimodal approaches for generic video summarization which take as input the full-length video and a textual description of its content and/or the video transcripts. The CLIP-It method of Narasimhan et al. (2021) [6] for query-based video summarization is also capable to produce generic video summaries using automatically-generated dense captions of the video content. In their subsequent work, Narasimhan et al. (2022) [7] described a transcript-based method for summarizing instructional videos by selecting important steps of the procedure that are most relevant to the task, as well as salient in the video, i.e., referenced in the speech. Zhong et al. (2022) [21] presented an unsupervised video summarization approach that learns how to estimate the frames’ importance and create semantically representative video summaries by minimizing the distance of learnable representations of the video content and its textual description, in a common latent space. He et al. (2023) [8] presented a unified multimodal transformer-based model which can align and attend multimodal inputs leveraging time correspondence (i.e., video and transcript), and selects the most importance frames of the video (to form a keyframe-based video summary) and the most importance sentences of the transcript (to produce a text summary). Finally, Argaw et al. (2024) [22] trained a multimodal method for video summarization using ground-truth pseudo-summaries obtained by prompting an LLM to extract the most critical and informative moments from ASR transcripts and mapping back the resulting textual summary to the relevant video segments. Their method creates a video summary by taking into account the visual content and a long-form description of it or the video transcripts.
The aforementioned multimodal approaches utilize additional information about the content of the full-length video. As a consequence, the generated summaries provide a synopsis of the entire video and they are not adapted to any specific requests about their content. Differently from these methods, SD-VSum takes into account information about the content of the desired summary, and thus it is capable to produce video summaries that do not contain redundant information and are customized to the users’ demands.
Table 1: Overview of the different datasets for text/query-driven video summarization from the literature. The original VideoXum dataset that is used as a basis for our S-VideoXum dataset, is included as a reference.
Dataset Domains Videos Duration
Annotations
per video
Type of annotations for text/
query-driven summarization
Data split
UTE [23] 1 4 3 - 5 hr. 3 textual summaries train, test
TVE [24] 1 4 45 min. 3 textual summaries train, test
QFVS [3] 1 4 3 - 5 hr. 46 query-based video summaries train, test
RAD [11] 22 190 2 - 3 min. 5 query-relevance labels train, validation, test
SumMe [25] - 25 1 - 6 min. 3-5 fragment-level descriptions train, test
TVSum [26] 10 50 1 - 11 min. 3-5 fragment-level descriptions train, test
ARS [17] 2 10 2 - 12 min. 10 query-based video summaries train, test
VideoXum [9] open domain 14,000 up to 12.5 min. 10 - train, validation, test
S-VideoXum (Ours) open domain 11,908 up to 12.5 min. 10 video summary descriptions train, validation, test
2.3 Datasets
An overview of the datasets that were used in the literature for text/query-driven video summarization is provided in Table 1. The earliest used ones by Sharghi et al. in [2], were the UT Egocentric (UTE) [23] and the TV Episodes (TVE) [24] datasets. UTE contains four egocentric videos $3-5$ hours long that were recorded in an uncontrolled environment from a first-person view wearable camera. TV Episodes includes four videos $45$ min. long that were captured from a third person’s viewpoint. The videos of the UTE and TV Episodes have been annotated by human subjects in [24] with one-sentence descriptions at the level of short video fragments that last $5$ and $10$ sec., respectively. Three reference text-based summaries were defined per video as a subset of the textual annotations, and used for evaluating the output of the query-driven summarization method based on the ROUGE-SU metric [27]. The UTE dataset was further extended by Sharghi et al. in [3], by adding human-annotated video summaries for $46$ user queries that were defined using two or three concepts per video, forming the Query-Focused Video Summarization (QFVS) dataset. Vasudevan et al. [11] introduced the RAD dataset which contains $190$ videos that were retrieved from YouTube using different queries and annotated with query relevance labels at the frame and fragment level. Labels range from 0 (Bad) to 3 (Very Good), where “Bad” indicates that the frame is both irrelevant and of low aesthetic quality (e.g. blurred, bad contrast, etc.). This dataset was extended by Huang et al. in [5], by computing relevance scores for each query-video pair of the dataset. The SumMe [25] and TVSum [26] benchmarking datasets for generic video summarization were extended by Wei et al. [4] with fragment-level descriptions of the visual content, to train a video summarization method that takes into account the video semantics. Finally, Cizmeciler et al. [17] introduced the Activity Related Summaries (ARS) dataset that contains $10$ YouTube videos of $2$-$12$ min. duration. These videos were annotated by humans at the shot level using a pre-defined list of $60$ attributes, and $10$ ground-truth summaries were defined for each video using one/two-word queries from a list of $142$ queries.
As shown in Table 1, contrary to the existing datasets, the extended S-VideoXum dataset for script-driven video summarization: i) contains a remarkably large number of videos from different domains with diverse visual content, ii) provides $10$ different ground-truth summaries per video, thus allowing to train a summarization method to produce different summaries for a given video based on the users’ preferences about the content of each summary, and iii) is divided in train, validation and test sets, enabling the selection of a well-trained model that generalizes well to unseen data.
3 Problem Statement and Dataset
3.1 Problem formulation
Let us assume an input video of $N$ frames and a user-provided script (a free-form text description) of $M$ sentences outlining the content of the desired video summary; different sentences of the script may refer to different parts of the full-length video with varying visual and semantic content. Then, script-driven video summarization aims to select a subset of frames or fragments of the video that are both relevant to one or more sentences of the user-provided script, and important for an overall understanding of the full-length video. The selected frames/fragments should form a video summary that does not exceed a predefined length $L$. Following the paradigm in the video summarization literature [1], $L$ is typically set to $15\%$ of the full-length video duration.
Another task that aims to spot parts of the video that are relevant to a freeform query is the task of temporal sentence grounding in videos [28] (a.k.a. video grounding). Nevertheless, this task tries to localize one target segment from the full-length video, that corresponds to a given sentence query. On the contrary, script-driven video summarization aims to generate a concise summary, and in this process it needs to select multiple video fragments that relate to the complex query, as well as other less relevant parts that are necessary for providing a synopsis of the full-length video.
3.2 Script-driven video summarization dataset
To overcome observed weaknesses of small video summarization datasets, related to the sensitivity of the evaluation outcomes to different data splits (as shown in [29]) and the inability to select a well-trained model in the absence of a validation set (discussed in [1]), we focused on the VideoXum large-scale dataset for cross-modal video summarization [9]. VideoXum includes $14,000$ open-domain videos up to $12.5$ min. long with diverse visual content, from the ActivityNet Captions dataset [16]. Each video is associated with multiple ground-truth video summaries - in the form of frame-level binary scores which denote the inclusion (label “1”) or not (label “0”) of a frame in the video summary - that were obtained by $10$ different human annotators. The dense video captions from the ActivityNet Captions dataset that are considered as text summaries in [9] provide a high-level description of the full-length video and thus are more appropriate for multimodal generic video summarization purposes. The existence of multiple ground-truth summaries per video ($10$ in total) makes the VideoXum dataset well-suited for the script-driven video summarization task, as it allows to train a summarization method to generate different summaries for a given video, driven by the provided script about the content of each summary. Furthermore, its significantly large size allows to split the available videos into training, validation and test set. Such a data split supports the extensive training of a model using a very wide set of training samples, the selection of a well-trained model that exhibits good generalization capacity through the use of an adequately big validation set, and the robust evaluation of its performance using a large number of testing samples.
To make VideoXum suitable for training and evaluation of script-driven video summarization methods, we extended it by producing natural language descriptions of the different ground-truth summaries that are available per video. For this, we employed the publicly-available [^1] https://huggingface.co/llava-hf/LLaVA-NeXT-Video-7B-hf state-of-the-art Large Multimodal Model LLaVA-NeXT-7B [10]. To align with the adopted video representation strategy in [9], we processed each ground-truth video summary using one frame per second and obtained the natural language descriptions using the following prompt: “Describe the important scenes in the video”. The maximum number of generated tokens was set equal to $200$, and a $4$-bit quantization was applied to reduce computational cost. Moreover, to allow performance comparisons with multimodal approaches for generic video summarization (which are compatible with our method in terms of the used data modalities), we produced natural language descriptions of the full-length videos, as well. Once again, we sampled these videos keeping one frame per second and generated descriptions using the LLaVA-NEXT-7B model and the same prompt.
Figure 2: Overview of the SD-VSum network architecture. Given an input video and a user script outlining the content of the desired summary, SD-VSum generates a summary conditioned on the script. Visual and text embeddings are obtained using the corresponding encoders of a pretrained multimodal vision and language model, and fused using a trainable cross-modal attention mechanism. The cross-modal embeddings in the output of this mechanism, that encode relationships between the different modalities and the position of the video frames, are forwarded to a trainable Transformer-based frame scoring mechanism, which computes frame-level importance scores. These scores are finally used by a frame/fragment selection component that forms the video summary based on a predefined temporal fragmentation of the video and a time-budget about the summary duration.
Since some of the videos from the VideoXum dataset were not publicly-available during this work, the extended S-VideoXum dataset for query-driven video summarization differs from the original one in terms of number of videos, including data for $11,908$ videos. However, a statistical analysis on the obtained videos showed that our dataset maintains the same distribution of training, validation and test samples with the original VideoXum dataset. In particular, we use $6,782$ samples for training, $3,419$ for validation and $1,707$ for testing purposes. The generated natural language descriptions for the ground-truth summaries and the full-length videos of the S-VideoXum dataset, along with the extracted (CLIP-based) embeddings from visual and textual data and the used data splits in our experiments, are publicly-available on GitHub at: https://github.com/IDT-ITI/SD-Vsum.
4 Proposed Approach
4.1 Network architecture
An overview of the SD-VSum network architecture is provided in Fig. 2. Assuming an input video of $N$ frames (following sampling), and a user script of $M$ sentences that outlines the content of the desired summary, SD-VSum produces a video summary conditioned on the script and a time-budget about its length. A set of visual embeddings ($\boldsymbol{X}=\{\boldsymbol{x_{n}}\}_{n=1}^{N}$) of size $D$ ($\boldsymbol{x_{n}}=\{x_{r}\}_{r=1}^{D}$) are obtained for the video frames, using the visual encoder of a pretrained multimodal embedding model. Similarly, a set of text embeddings ($\boldsymbol{Y}=\{\boldsymbol{y_{m}}\}_{m=1}^{M}$) of the same size ($\boldsymbol{y_{m}}=\{y_{r}\}_{r=1}^{D}$) are computed for the different sentences of the user script (one per sentence), using the text encoder of the aforementioned model. These embeddings are then given as input to a multi-head cross-modal attention mechanism, as shown in Fig. 3. Given the $h^{th}$ attention head of this mechanism, the visual embeddings $\boldsymbol{X}$ pass through a linear layer of size $D/H$, where $H$ denotes the number of heads, forming the Query ($\boldsymbol{Q}_{h}=\{\boldsymbol{q}_{n}\}_{n=1}^{N}$) matrix. The text embeddings $\boldsymbol{Y}$ pass through two different linear layers of size $D/H$, creating the Key ($\boldsymbol{K}_{h}=\{\boldsymbol{k}_{m}\}_{m=1}^{M}$) and Value ($\boldsymbol{V}_{h}=\{\boldsymbol{v}_{m}\}_{m=1}^{M}$) matrices. The Query and Key matrices undergo a matrix multiplication process that is followed by a softmax layer, formulating the values of the attention matrix ($\boldsymbol{A}_{h}=softmax(\boldsymbol{Q}_{h}\times\boldsymbol{K}_{h}^{T})\ and\ \boldsymbol{A}_{h}=\{a_{i,j}\}_{i=1,N}^{j=1,M}$). This matrix goes through a dropout layer and the outcome is multiplied with the Value matrix, defining the output of the $h^{th}$ attention head ($\boldsymbol{Z}_{h}=\{\boldsymbol{z_{n}}\}_{n=1}^{N}\ and\ \boldsymbol{z_{n}}=\{z_{r}\}_{r=1}^{D/H}$). The obtained representations in the output of all the attention heads are then concatenated, and the resulting representations pass through a linear layer of size $D$ and a positional encoding mechanism (similar to the one in [30]), formulating the overall cross-modal embeddings ($\boldsymbol{Z}=\{\boldsymbol{z}_{n}\}_{n=1}^{N}$) in the output of the attention mechanism. Subsequently, as depicted in Fig. 2, these embeddings undergo a dropout and normalization process and then are given as input to a trainable Transformer-based scoring mechanism which computes frame-level importance scores ($\boldsymbol{f}=\{f_{n}\}_{n=1}^{N}$). These scores are finally used by a frame/fragment selection component that forms the video summary based on a predefined temporal fragmentation of the video and a time-budget about the summary duration.
Figure 3: The processing pipeline of the utilized multi-head cross-modal attention mechanism. The visual embeddings form the Query matrix and the text embeddings form the Key and Value matrices based on trainable linear layers. The Query and Key are then multiplied and the computed values pass through a softmax layer, forming the Attention matrix. This matrix is multiplied with the Value, producing the output of an attention head. The outputs of all attention heads are concatenated, and the obtained representations pass through a linear layer and a positional encoding mechanism, formulating the overall cross-modal embeddings in the output of the attention mechanism.
So, in terms of network architecture, SD-VSum is most closely related with the CLIP-It architecture [6] for query-based and multimodal generic video summarization. CLIP-It uses a language-guided multi-head attention mechanism to fuse information from the visual and text modalities, and employs a Transformer-based scorer for assigning importance scores to the video frames. With respect to the representation of the input data, the video is represented using similar (CLIP-based) frame-level embeddings, but the query (in the case of query-based video summarization) or the dense video captions (in the case of multimodal generic video summarization) is represented using a condensed single-vector embedding. The latter is produced by a trainable linear layer that gets as input a vector formed by concatenating a fixed number of sampled (CLIP-based) embeddings of the input text, and projects it into a more condensed representation (of dimension $1\times 512$). As a consequence, the language-guided attention of CLIP-It encodes the dependence between the video frames and the user query in a more compact, and thus less informative, manner. On the contrary, the cross-modal attention mechanism of the proposed SD-VSum network architecture employs a more detailed representation of the input text at the sentence level, allowing SD-VSum to encode the frames’ dependence with different parts of the user script in a more effective way (using an $N\times M$ attention matrix, as shown in Fig. 3, rather than a more condensed $N\times 1$ matrix, as in CLIP-It). This main distinction leads to a noticeably advanced summarization performance, as documented in Section 5.3.
4.2 Training regime
Training of SD-VSum is performed in a supervised learning setting, comparing its output for a given pair of full-length video and user script, with the associated ground-truth summary (i.e., the one used to generate this particular script). In particular, we compute the Binary Cross Entropy (BCE) between the computed scores for the frames of the input video by SD-VSum ($\boldsymbol{f}=\{f_{n}\}_{n=1}^{N}$), and the assigned binary labels to these frames by a human annotator. The computed training loss is then back-propagated to compute the gradients and update all the different trainable parts of the network architecture. The use of the multiple available ground-truth summaries for each video of the dataset as individual training samples, practically increases tenfold the amount of training, validation and test samples, reported in Section 3.2.
A different training regime is followed when evaluating the performance of SD-VSum on the task of multimodal generic video summarization. In this case, the computed frame-level scores for a given pair of full-length video and textual description of its content, are compared with a single ground-truth summary that is formed by averaging the multiple human-annotated summaries for this video, at the frame-level. So, instead of comparing the output of SD-VSum with multiple different ground-truth annotations for the input video using BCE (a contradictory approach that could prevent the training of the network architecture), we calculate a single loss value by using an average ground-truth summary and the Mean Squared Error (MSE).
5 Experiments
5.1 Evaluation protocol
Following the evaluation approach in [9], we generate the video summary for a given input video and user script, by selecting the top-15% scoring frames by a video summarization model. Then, we quantify the similarity between the machine-generated and the ground-truth summary using the F-Score (as percentage). So, a given test video is matched with each one of the multiple available user scripts for it, and each one of the generated summaries is compared with the corresponding ground-truth summary. Through this process, we compute an F-Score for each pair of compared summaries and we average these scores to form the final F-Score for this video. After performing this for all test videos, we calculate the mean of the obtained F-Score values and result in an average score that indicates the method’s performance on the test set.
When evaluating the performance on multimodal generic video summarization we also use the evaluation protocol from [31]. In particular, we measure the alignment between the computed frame-level importance scores for a given video by a video summarization model, and the ground-truth frame-level importance scores for this video (obtained by averaging its multiple available binary ground-truth summaries, at the frame-level), using the Kendall’s $\tau$ [32] and Spearman’s $\rho$ [33] rank correlation coefficients. The computed $\tau$ and $\rho$ values for all test videos are then averaged, defining the performance of the video summarization model on the test set.
The selection of a well-trained model for evaluation on the videos of the test is made based on the recorded performance on the videos of the validation set. More specifically, after each training epoch we measure the performance of the trained model on the videos of the validation set. When then training process is completed, we keep the model with the highest validation-set performance in terms of F-Score, and assess its performance on the test set using the evaluation protocols described above.
5.2 Implementation details
Similarly to [9], videos are represented using one frame per second, and frame-level embeddings of the visual content (of size $D=512$) are obtained using a fine-tuned CLIP model on the data of VideoXum, that has been released by the authors of [9] [^2] Publicly-available at: https://videoxum.github.io/. The same model is used to extract sentence-level embeddings (of the same size $D=512$) from the generated scripts. With respect to the SD-VSum network architecture, the number of attention heads is set equal to $8$, and the Transformer-based frame scoring mechanism consists of a Transformer network [34], a linear layer with $512$ neurons, and a sigmoid function that computes the frame-level importance scores. The learning rate, dropout rate and L2 regularization factor are equal to $5\cdot 10^{-5}$, $0.5$ and $10^{-4}$, respectively. For the network’s weights initialization, we use the Xavier uniform initialization approach with gain = $\sqrt{2}$ and biases = $0.1$. Training is performed in a batch mode with a batch size equal to $4$ using the Adam optimizer, and stops after $50$ epochs. All experiments were carried out on a PC with an NVIDIA RTX 3090 GPU. To allow the reproduction of our results, the used data and the PyTorch implementation of SD-VSum will be made publicly-available upon paper’s acceptance.
5.3 Experimental comparisons and ablations
Table 2: Script-driven video summarization: performance comparison (F-Score (%)) with the CLIP-It method and several variants of the proposed SD-VSum network architecture.
Text
representation
# Heads
Scaling in
attention
F-Score
SD-VSum multiple vectors 8 No 24.8
CLIP-It [6] single vector 4 Yes 22.8
Variant 1 single vector 8 No 22.8
Variant 2 single vector 8 Yes 22.7
Variant 3 single vector 4 Yes 22.2
Variant 4 multiple vectors 8 Yes 24.2
The proposed SD-VSum network architecture was compared with CLIP-It [6] that exhibits state-of-the-art performance on query-driven and multimodal (generic) video summarization. The results in the upper part of Table 2 show that SD-VSum performs clearly better than CLIP-It on the extended S-VideoXum dataset, achieving an F-Score that is almost $2\%$ higher than the F-Score of CLIP-It. Since SD-VSum and CLIP-It are similar in terms of network architecture (as discussed in Section 4.1), we tried to spot what causes this noticeable difference in the script-driven summarization performance of these architectures. For this, we conducted an ablation study that involved variants of SD-VSum that resemble the network architecture of CLIP-It, as well as a variant that focuses on the use of a scaling process when computing the attention weights. In particular, we considered the following variants:
  • Variant #1 applies the same single vector representation approach for the user script, with CLIP-It.
  • Variant #2 extends the previous variant by containing the same number of attention heads with CLIP-It.
  • Variant #3 extends the previous variant by performing a scaling process when computing the attention weights (by dividing with $\sqrt{D}$, where $D$ is the size of the used embeddings), simulating further the language-guided attention mechanism of CLIP-It.
  • Variant #4 uses multiple embeddings for the user script and maintains the same number of attention heads ($8$), but performs the aforementioned scaling process when computing the attention weights.
The findings of our ablation study are reported in the lower part of Table 2. The performance of Variant #1 highlights the impact of producing and using a significantly more condensed representation for the user script, which leads to a noticeable drop in the script-driven video summarization performance (almost $2\%$ lower than SD-VSum). The performance of Variant #2 indicates the slight effect of the scaling process when using a single vector representation for the user script ($0.1\%$ lower than the performance of Variant #1), an effect that is more pronounced when the script is represented using multiple embeddings (as denoted by the performance of Variant #4 which is $0.6\%$ lower than that of SD-VSum). Finally, the reduction of the number of employed attention heads leads to further performance reduction ($0.5\%$ lower than the performance of Variant #2). In general, the observed performance for the CLIP-It method and the variants of SD-VSum that utilize the same approach for representing the user script, documents the contribution of using a more coarse-grained sentence-level representation of the user script and points out the effectiveness of the integrated cross-modal attention mechanism in the SD-VSum network architecture.
Table 3: Performance comparison between the proposed SD-VSum network architecture for script-driven video summarization and literature networks for visual-based and multimodal video summarization, in terms of F-Score (%) and Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients.
F-Score $\tau$ $\rho$
Script-driven
summarization
SD-VSum 24.8 N/A N/A
Multimodal summarization SD-VSum 23.6 0.188 0.247
VTSUM-BLIP [9] 23.5 0.196 0.258
CLIP-It [6] 22.8 0.154 0.203
Visual-based summarization VSUM-BLIP [9] 23.1 0.185 0.246
PGL-SUM [30] 22.0 0.153 0.203
Our next experiments focused on assessing the capacity of generic video summarization methods to tackle the needs of the script-based video summarization task. For this, we evaluated the performance of SD-VSum and CLIP-It after using as input a natural language description of the full-length video (instead of the desired video summary). Moreover, we ran experiments with a state-of-the-art method for visual-based video summarization (PGL-SUM [30]) that estimates the frames’ importance using multi-head attention mechanisms that model the frames’ dependence at different levels of granularity. As a note, the computation of the Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients in the case of script-driven video summarization is not possible, since the output of the network architecture is compared with a binary ground-truth summary. Our comparisons included also the visual-based and multimodal network architectures that were evaluated by the authors of the VideoXum dataset, after taking into account the scores reported in [9]. The outcomes of this investigation, presented in Table 3, show that the visual-based video summarization approaches are the least effective ones; thus, using a description of the full-length video as auxiliary input information, seems to be beneficial for the summarization process. Concerning the multimodal approaches, the performance of SD-VSum is clearly better than the one of CLIP-It, and highly competitive to the one of VTSUM-BLIP. Nevertheless, it is lower than the exhibited performance when SD-VSum is used for script-driven video summarization. These findings indicate the importance of script-driven video summarization technologies, and their competency to produce video summaries that are better tailored to the needs of different users.
Finally, in terms of computational complexity, we counted the time needed per training epoch and the number of learnable parameters. The proposed SD-VSum network architecture is a bit less complex than CLIP-It, due to the lack of a trainable linear layer that is responsible for producing the single vector representation of the input text. This difference is reflected in Table 4, by the slightly reduced time needed per training epoch and the smaller number of trainable parameters of SD-VSum. So, the observed improvements in the summarization performance are gained through a less memory- and computational-demanding training process.
Table 4: Comparison in terms of computational complexity.
Network
architecture
Training time
(min. / epoch)
# Parameters
(in Millions)
SD-VSum 25.25 45.19
CLIP-It [6] 26.17 45.98
5.4 Qualitative analysis
Our qualitative analysis aimed to investigate the extend to which the output of script-driven video summarization is more tailored to the needs of different users, compared to the generic video summaries that are produced by multimodal approaches. For this, we compared the overlap (expressed in F-Score (%)) between the multiple available ground-truth summaries for $50$ videos of the extended S-VideoXum dataset, and the produced summaries when the proposed SD-VSum is used for script-driven and multimodal video summarization, respectively. The pool of videos was defined by choosing the ones that showed the major diversity in the computed overlap when comparing the multimodal-based video summary with each different ground-truth summary, and thus were the most challenging ones for a multimodal video summarization approach. The obtained scores, that are illustrated in the form of heatmap values in Fig. 4, demonstrate that the output of multimodal video summarization (see the heatmap on the right side) shows the same weak overlap with the summaries of the different human annotators, and thus is poorly adapted to different needs about the content of the summary. On the contrary, the output of script-driven video summarization is clearly more aligned to the preferences of each different annotator, for most of the selected videos. These findings justify once again the motivation behind the development and training of network architectures for script-driven video summarization, and highlight the capacity of the proposed SD-VSum network architecture to produce video summaries that are well-tailored to each user’s needs about their content.
Figure 4: The overlap (expressed in F-Score (%)) between the multiple available ground-truth summaries for $50$ videos of the extended S-VideoXum dataset, and the produced video summaries by SD-VSum when it is used for script-driven (left side) and multimodal video summarization (right side).
6 Conclusions
In this paper, we introduced a new task that aims to create a video summary of the full-length video by choosing the parts showing the highest relevance with a user-provided script that outlines the content of the desired summary. To support our research, we extended the VideoXum large-scale dataset for video summarization by generating natural language descriptions of the different ground-truth summaries that are available per video. Using the extended S-VideoXum dataset, we trained and evaluated a new network architecture for script-driven video summarization (SD-VSum), that integrates a cross-modal attention mechanism to align and fuse information from the different data modalities. Our quantitative evaluations highlighted the advanced performance of SD-VSum compared to state-of-the-art methods for query-driven and generic (unimodal and multimodal) summarization from the literature, and pointed out the impact of using coarse-grained representations for the input text. Our qualitative analysis demonstrated the capacity of SD-VSum to incorporate knowledge about the user’s needs and produce well-customized video summaries.